# Towards Scalable Automated Alignment of LLMs: A Survey

### v2

@@ 1 Introduction
Recent years have witnessed the rapid advancements of large language models (LLMs), which have dramatically reshaped the landscape of artificial intelligence (Ouyang et al., 2022; Touvron et al., 2023; OpenAI, 2023c). Alignment is at the core of shaping behaviors of LLMs corresponding to human intentions and values (Yao et al., 2023a; Shen et al., 2023b), e.g., teaching LLMs to follow “helpful, harmless and honest (HHH)” principles during responding (Askell et al., 2021). As a result, increasing efforts have been made for aligning LLMs to meet the human requirements, which makes it a hotspot research direction in LLM era (Wang et al., 2023g, 2024g; Ji et al., 2023).
@@
To address these challenges, automated alignment has drawn great attention very recently (Yuan et al., 2024b; Chen et al., 2024g). Unlike previous methods that relied on human annotation to obtain alignment signals, the goal of automated alignment is constructing scalable and high-quality alignment systems with minimal human intervention. Therefore, automated alignment has the potential to address the core challenges posed by the rapid development of LLMs, where human annotation is either infeasible or extremely expensive. For automated alignment, the most crucial part is to find a scalable alignment signal that can replace human manually-created preference signals and remain effective amid the rapid development of LLMs.
@@
The rest of this survey is organized as follows: Section 2 describes the scope of automated alignment covered in this survey, as well as the our taxonomy. Section 3-6 provide a detailed introduction to the progress and limitations of the four aforementioned representative directions in automated alignment. Section 7 explores the underlying mechanisms of automated alignment. And we include a overall conclusion of this survey in Section 8111We openly released a corresponding paper list which will be regularly updated on https://github.com/cascip/awesome-auto-alignment.
@@
Bias §3, for tree=acquisition [From features of LLMs, [Uncertainty Filtering, acquisition [Self-Consistency (Wang et al., 2023e); Self-Improve (Huang et al., 2023a); West-of-N (Pace et al., 2024); etc., acquisition_work] ] [Self-Judge/Critic/Refine, acquisition [ Constitutional AI (Bai et al., 2022b); Tree-of-Thought (Yao et al., 2023b); Self-Rewarding (Yuan et al., 2024b); etc.,acquisition_work ] ] [Context Distillation, acquisition [Lab for Alignment (Askell et al., 2021); Dromedary (Sun et al., 2023d); Llama-2-Chat (Touvron et al., 2023); RLCD (Yang et al., 2024b); etc., acquisition_work] ] ] [From organization of LLMs, [Task Decomposition, acquisition [Least-to-Most (Zhou et al., 2023b); IDA (Christiano et al., 2018); etc., acquisition_work] ] [Self-play, acquisition [SPIN (Chen et al., 2024g); Consensus Game (Jacob et al., 2024); etc., acquisition_work] [Debate (Irving et al., 2018); SPAG (Cheng et al., 2024); etc., acquisition_work] ] ] ] [Behavior
@@
Feedback §5, for tree=probing [Scalar Reward, [RLHF , probing [ InstructGPT (Ouyang et al., 2022); DPRM (Li et al., 2024a); etc., probing_work] ] [RLAIF, probing [RLAIF (Lee et al., 2023); RLCD (Yang et al., 2024b); etc., probing_work] ] [Feedback-guided Decoding, probing [Critic-driven Decoding (Lango and Dusek, 2023); RAD (Deng and Raffel, 2023); etc., probing_work] ] [Filtering SFT Data, probing [Quark (Lu et al., 2022); RRHF (Yuan et al., 2023a); RAFT (Dong et al., 2023); etc., probing_work] ] ] [Binary Verifier, [Outcome Verifier , probing [V-STaR (Hosseini et al., 2024); SORMs (Havrilla et al., 2024); etc., probing_work] ] [Process Verifier, probing [MATH-SHEPHERD (Wang et al., 2024e); MiPS (Wang et al., 2024h) etc., probing_work] ] ] [Text Critic, probing [ILF (Scheurer et al., 2022); LEMA (An et al., 2024); etc., cus_probing_work] ] ] [Environment
@@
Mechanism, application [LIMA (Zhou et al., 2023a); Rethinking (Ren et al., 2024); URIAL (Lin et al., 2024a); ICL&IT (Duan et al., 2023); Behavior shift (Wu et al., 2023a) etc., application_work] ] [Inner Workings of Self-feedback,application [GV-consistency (Li et al., 2024g); CriticBench (Lin et al., 2024b); Self-Rewarding (Yuan et al., 2024b); Humback (Li et al., 2024f); LLM-as-a-Judge (Zheng et al., 2024b) etc., application_work] ] [Feasibility of Weak-to-strong, application [Easy2Hard (Sun et al., 2024b; Hase et al., 2024); Weak2Strong (Burns et al., 2023); Principle2Behavior (Bai et al., 2022b; Sun et al., 2023d) etc., application_work] ] ] ]
@@ 3 Aligning Through Inductive Bias
    It does not require additional supervisory signals beyond the model itself, thus avoiding the high cost of obtaining additional annotated data. This is particularly relevant given the current scenario where training data is becoming scarce or has already been exhausted  [^2] Within the context of alignment discussed in this paper, we expect the model to continuously improve its helpfulness, thereby providing more effective assistance for alignment process. The scope of alignment actually represents a post-pretraining process rather than steering the models. (Xue et al., 2023).
@@
After conducting a thorough review of the relevant literature, we find that current efforts towards self-improvement solely through language model itself can be decomposed into a set $\mathcal{H}$ of five inductive biases. These inductive biases fall into two broad categories: 1) Those stemming from inherent features of LLMs (§3.1), and 2) Those arising from the organization of LLMs (§3.2).
For each type of inductive biases, we will begin by introducing its origin. Following this, we will enumerate the works that employ this inductive bias as a single-step policy improvement strategy. Next, we will discuss the works that iteratively train with it, aiming for continuous improvement. Finally, we will address open research problems associated with the given inductive bias.
@@ 3.1.1 H1: Uncertainty as a indicator of helpfulness
In the machine learning literature, early applications of this inductive bias were evident in a series of works using self-training (Scudder, 1965) for semi-supervised learning (Nigam and Ghani, 2000; Amini and Gallinari, 2002). The basic paradigm of these works involves using learners trained on supervised data to continue learning from confidently classified unlabeled data, thereby enhancing supervised learning performance with unlabeled data. This approach has been applied in classification tasks with methods such as Pseudo-label (Lee et al., 2013; Ferreira et al., 2023) and Entropy Minimization (Grandvalet and Bengio, 2004). He et al. (2020) extended this approach to sequence generation NLP tasks, highlighting that biased sampling and noise perturbation are key factors for the success of self-training in these tasks. Pace et al. (2024) extended the paradigm of self-training to the alignment problem, improving the robustness of reward models by allowing them to iteratively learn from the highest-scoring and lowest-scoring answers in the candidate pool generated for a query.
@@
To solidify such enhancement, Self-Improve (Huang et al., 2023a) views CoT with Self-Consistency as a policy improvement operator, substantially improving the reasoning capabilities and potential of LLMs through iterative learning of the reasoning paths obtained from Self-Consistency. Zhang and Parkes (2023) demonstrate that LMs could self-improve on large number addition problems by curriculum “distilling” the CoT answers to the direct answers without explicit reasoning. Quiet-STAR (Zelikman et al., 2024) considers the advantage of the influence of rollout rationale on the probabilities of subsequent tokens as a feedback signal, encouraging the model to generate a more helpful implicit thought process using reinforcement learning techniques.
@@ 3.1.2 H2: LLMs can judge, critic, refine and more
Judge refers to determining the quality of model responses. The judgment standards are typically incorporated into instructions as principles or guidelines, enabling regulators to oversee LLM behavior in a more scalable manner (Bai et al., 2022b; Yuan et al., 2024b) compared to heavy reliance on human annotators for feedback (Bai et al., 2022a). This approach allows for timely regulation, aiding in the flexible and controlled alignment process of language models, which can help prevent issues like reward hacking during iterative training (Sun et al., 2023c) and facilitate on-policy reinforcement learning training (Guo et al., 2024a).
Self-judging can manifest in two primary forms: 1) Differentiating the relative quality of two responses (AI Feedback (Bai et al., 2022b)), resulting in evaluation outcomes represented as a partial order. For instance, Tan et al. (2023) employ prompts to compare which answer better adheres to the HHH principles. They then distill the chosen option back into the model to further enhance its judging capability. Bai et al. (2022b) prompted the model to select superior responses based on sampled principles and subsequently employed a preference process to achieve a Pareto improvement of the model. 2) Providing an absolute score for a response (LLM-as-a-judge (Zheng et al., 2024b)), with the evaluation results in scalar form. Yao et al. (2023b), Besta et al. (2024) and Xie et al. (2023) introduce real-time evaluation modules for thought states during reasoning process. These modules serves as a prior during the search process, assisting the model in exploring the action space for questions requiring deliberate thinking. Similarly, RAIN (Li et al., 2024h) utilizes a binary scoring prompt for self-evaluating whether the generation is likely to be harmful, thereby enhancing response safety combined with an inference-time tree search. Yuan et al. (2024b) employ a five-point judge prompt to score the model’s instruction-response outputs, then converted the scores into a partial order for iterative training using DPO.
@@
Critique refers to generating modification suggestions. By leveraging LLM itself for critique, the suggestions can address errors and deficiencies, such as mistakes in summarization (Saunders et al., 2022), machine translation (Fernandes et al., 2023), math reasoning (Lin et al., 2024b), decision-making and programming tasks (Saunders et al., 2022). The suggestions can also pertain to abstract value criteria, such as the HHH related principles (Chen et al., 2024e; Bai et al., 2022b).
Refine refers to the ability that LLMs can improve the given text. Most of the work on self-refine is based on natural language reasons provided by the critic module to modify the response (Bai et al., 2022b; Tan et al., 2023; Madaan et al., 2023). Some research also demonstrates the possibility of making modifications directly based on scalar rewards (Shinn et al., 2023). The less informative critic can be more challenging for LLMs since they must complete more information by themselves through reasoning. Another line of work uses LLMs to refine the prompts themselves (Fernando et al., 2023; Yang et al., 2023a).
@@
    Conducting causal studies on the formation process of models’ judging, critiquing, and refining abilities, and investigating what forms of pre-training and fine-tuning data can influence these abilities.
@@ 3.1.3 H3: LLMs can effectively learn in-context
However, prepending few-shot exemplars in the context above can make inference inefficient (Gim et al., 2023) and interfered with the unrelated queries (Shi et al., 2023). Therefore, self-generated labels obtained from ICL could be directly used as pseudo labels, and distilled back into the LLMs only paired with the query. This paradigm is known as Context Distillation (Askell et al., 2021). For instance, in the alignment process of Llama-2 (Touvron et al., 2023), Context Distillation is used to alleviate the problem of long-term dependencies of system prompts. In Dromedary (Sun et al., 2023d), the base language model is transformed into a safe and helpful aligned model with minimal annotations by directly training on samples acquired from multiple ICL processes. Padmanabhan et al. (2023) demonstrate that Context Distillation can also be used to inject new knowledge to models by learning continuations from entity definitions. Furthermore, Yang et al. (2024b) illustrate the effectiveness of distilling the preference pair generated by contrastive in-context constraints back into the model.
@@ 3.2.2 H5: Self-play
As discussed in H2, the judge and critic model is commonly considered a type of discriminator. For example, Yuan et al. (2024b) utilized rewards from LLM-as-a-Judge to identify high-quality and low-quality responses from the generator, optimizing the generator towards the higher quality ones. However, the adversarial setting between the discriminator and the generator is limited because the only assumption is that the discriminator’s ability can be improved with general helpful training. The discriminator remains almost static (the prompts are unchanged) during training, making it possible for the generator to be over-optimized against the discriminator, leading to reward hacking. Effectively improving the judge and critic module alongside the generator is thus a crucial problem. One reasonable approach is to use the online AIF setting, as suggested by Guo et al. (2024a). Moreover, Cheng et al. (2023) propose an adversarial game between the policy and reward model, facilitated through a constrained min-max loss. Another way to introduce a more adversarial setting is to optimize the gaming problem at inference time, as demonstrated in the Consensus Game by Jacob et al. (2024). It employs the piKL no-regret learning algorithm to iteratively update the strategies of both the generator and discriminator, converging to a Nash equilibrium. This equilibrium strategy is then used to rank candidate responses, prioritizing those agreed upon by both players.
@@
The debate paradigm (Irving et al., 2018) is largely inspired by factored cognition and AlphaGo (Silver et al., 2018). In the learning algorithm of AlphaGo, three distinct components are integrated: a player, a counterpart (itself), and a value model that evaluates the win rate associated with each board state. By employing Monte Carlo Tree Search (MCTS), the algorithm conducts rollouts, which are simulated self-play sequences that extend until a game’s conclusion. These rollouts enhance the accuracy of the value estimates through backward updates based on the outcomes, concurrently refining the policy by capitalizing on strategies that have previously led to victories.
@@
The game of Go shares similarities with tackling the scalable oversight problem using natural language debate. In the beginning or middle of a Go game, even experienced experts may find it difficult to judge which side has a higher probability of winning, just as humans have a small probability of judging problems that exceed human knowledge levels. However, as the game nears its end, the outcome usually becomes clear that even a non-expert judge can confidently evaluate the board generated by the Go masters. For a debate competition, the winner can typically be summarized by the judges.
@@
This provide a possible oversight solution to build trustworthy superhuman AI systems. Irving et al. (2018) show that honesty is better strategy than lie in debate paradigm through proof-of-concept experiments. As an extension of this, Brown-Cohen et al. (2023) propose a new set of debate protocols, wherein the honest strategy can always succeed through a simulation involving only a polynomial number of steps. Khan et al. (2024) conduct a thorough empirical study on the feasibility of implementing the debate paradigm on LLMs: it was found that the debate paradigm can significantly enhance truthfulness, and more persuasive (Anthropic, 2024a) debaters lead to more truthful outcomes.
@@
Apart from the classic natural language debate, a series of works have implemented the debate paradigm in game scenarios. A representative arena is the bargaining task (Nash et al., 1950). Fu et al. (2023a) focus on zero-sum variants of bargaining, where the balloon seller aims to sell at a higher price while the buyer seeks a lower price. They observed significant variations in bargaining capabilities among different LLMs and their capacity to learn from play experiences and feedback. Cheng et al. (2024) implement the adversarial language game adversarial taboo (Yao et al., 2021), where an attacker and a defender engage in a conversation centered around a target word visible only to the attacker. The attacker subtly induces the defender to unconsciously utter the target word, while the defender tries to avoid doing so and guess the word from the context. Both players acquire basic gaming skills through imitation learning from a teacher LLM and then refine their strategies through self-play. Interestingly, less capable player LLMs not only improve their win rates in this specific game but also enhance their general reasoning abilities. Ma et al. (2023a) introduce the red-teaming game, a more intricate adversarial team game where LLMs are initialized as a joint set of red-teaming policies to prompt the target LLM to produce harmful content. They propose a solver to ensure the final meta-strategy approximates a Nash equilibrium within a certain $\epsilon$ margin. Zheng et al. (2024c) suggests addressing the alignment problem by allowing an attacker to prompt a defender LLM to generate answers that might result in low rewards, while the defender tries to maximize the rewards of these prompts. The solution of this game is considered an iterative min-max optimization process with constraints.

### v3

@@ 1 Introduction
Recent years have witnessed the rapid advancements of large language models (LLMs), which have dramatically reshaped the landscape of artificial intelligence (Ouyang et al., 2022; Touvron et al., 2023; OpenAI, 2023c). Alignment is at the core of shaping behaviors of LLMs corresponding to human intentions and values (Yao et al., 2023a; Shen et al., 2023b), e.g., teaching LLMs to follow “helpful, harmless and honest (HHH)” principles during responding (Askell et al., 2021). As a result, increasing efforts have been made for aligning LLMs to meet the human requirements, which makes it a hotspot research direction in LLM era (Wang et al., 2023g, 2024h; Ji et al., 2023).
@@
To address these challenges, automated alignment has drawn great attention very recently (Yuan et al., 2024c; Chen et al., 2024g). Unlike previous methods that relied on human annotation to obtain alignment signals, the goal of automated alignment is constructing scalable and high-quality alignment systems with minimal human intervention. Therefore, automated alignment has the potential to address the core challenges posed by the rapid development of LLMs, where human annotation is either infeasible or extremely expensive. For automated alignment, the most crucial part is to find a scalable alignment signal that can replace human manually-created preference signals and remain effective amid the rapid development of LLMs.
@@
The rest of this survey is organized as follows: Section 2 describes the scope of automated alignment covered in this survey, as well as the our taxonomy. Section 3-6 provide a detailed introduction to the progress and limitations of the four aforementioned representative directions in automated alignment. Section 7 explores the underlying mechanisms of automated alignment. And we include a overall conclusion of this survey in Section 8.
@@
Bias §3, for tree=acquisition [From features of LLMs, [Uncertainty Filtering, acquisition [Self-Consistency (Wang et al., 2023e); Self-Improve (Huang et al., 2023a); West-of-N (Pace et al., 2024); etc., acquisition_work] ] [Self-Judge/Critic/Refine, acquisition [ Constitutional AI (Bai et al., 2022b); Tree-of-Thought (Yao et al., 2023b); Self-Rewarding (Yuan et al., 2024c); etc.,acquisition_work ] ] [Context Distillation, acquisition [Lab for Alignment (Askell et al., 2021); Dromedary (Sun et al., 2023d); Llama-2-Chat (Touvron et al., 2023); RLCD (Yang et al., 2024b); etc., acquisition_work] ] ] [From organization of LLMs, [Task Decomposition, acquisition [Least-to-Most (Zhou et al., 2023b); IDA (Christiano et al., 2018); etc., acquisition_work] ] [Self-play, acquisition [SPIN (Chen et al., 2024g); Consensus Game (Jacob et al., 2024); etc., acquisition_work] [Debate (Irving et al., 2018); SPAG (Cheng et al., 2024); etc., acquisition_work] ] ] ] [Behavior
@@
Feedback §5, for tree=probing [Scalar Reward, [RLHF , probing [ InstructGPT (Ouyang et al., 2022); DPRM (Li et al., 2024a); etc., probing_work] ] [RLAIF, probing [RLAIF (Lee et al., 2023); RLCD (Yang et al., 2024b); etc., probing_work] ] [Feedback-guided Decoding, probing [Critic-driven Decoding (Lango and Dusek, 2023); RAD (Deng and Raffel, 2023); etc., probing_work] ] [Filtering SFT Data, probing [Quark (Lu et al., 2022); RRHF (Yuan et al., 2023a); RAFT (Dong et al., 2023); etc., probing_work] ] ] [Binary Verifier, [Outcome Verifier , probing [V-STaR (Hosseini et al., 2024); SORMs (Havrilla et al., 2024); etc., probing_work] ] [Process Verifier, probing [MATH-SHEPHERD (Wang et al., 2024e); MiPS (Wang et al., 2024i) etc., probing_work] ] ] [Text Critic, probing [ILF (Scheurer et al., 2022); LEMA (An et al., 2024); etc., cus_probing_work] ] ] [Environment
@@
Mechanism, application [LIMA (Zhou et al., 2023a); Rethinking (Ren et al., 2024); URIAL (Lin et al., 2024a); ICL&IT (Duan et al., 2023); Behavior shift (Wu et al., 2023a) etc., application_work] ] [Inner Workings of Self-feedback,application [GV-consistency (Li et al., 2024g); CriticBench (Lin et al., 2024b); Self-Rewarding (Yuan et al., 2024c); Humback (Li et al., 2024f); LLM-as-a-Judge (Zheng et al., 2024b) etc., application_work] ] [Feasibility of Weak-to-strong, application [Easy2Hard (Sun et al., 2024b; Hase et al., 2024); Weak2Strong (Burns et al., 2023); Principle2Behavior (Bai et al., 2022b; Sun et al., 2023d) etc., application_work] ] ] ]
@@ 3 Aligning Through Inductive Bias
    It does not require additional supervisory signals beyond the model itself, thus avoiding the high cost of obtaining additional annotated data. This is particularly relevant given the current scenario where training data is becoming scarce or has already been exhausted  [^1] Within the context of alignment discussed in this paper, we expect the model to continuously improve its helpfulness, thereby providing more effective assistance for alignment process. The scope of alignment actually represents the post training process rather than steering the models. (Xue et al., 2023).
@@
After conducting a thorough review of the relevant literature, we find that current efforts towards self-improvement solely through language model itself can be decomposed into a set $\mathcal{H}$ of five inductive biases. These inductive biases fall into two broad categories: 1) Those stemming from inherent features of LLMs (§3.1), and 2) Those arising from the organization of LLMs (§3.2). The objective of each of these inductive biases can be summarized as a simple rule: Heuristically transform test-time computation into the model’s alignment. (Snell et al., 2024) The additional computation of each inductive bias is depicted in the shaded regions of the illustrative figures.
For each type of inductive biases, we will begin by introducing its origin. Following this, we will enumerate the works that employ this inductive bias as a single-step policy improvement operator. Next, we will discuss the works that iteratively train with it, aiming for continuous improvement. Finally, we will address open research problems associated with the given inductive bias.
@@ 3.1.1 H1: Uncertainty as a indicator of helpfulness
In the machine learning literature, early applications of this inductive bias were evident in a series of works using self-training (Scudder, 1965) for semi-supervised learning (Nigam and Ghani, 2000; Amini and Gallinari, 2002). The basic paradigm of these works involves using learners trained on labeled data to continue learning from confidently classified unlabeled data, thereby enhancing supervised learning performance with unlabeled data. This approach has been applied in classification tasks with methods such as Pseudo-label (Lee et al., 2013; Ferreira et al., 2023) and Entropy Minimization (Grandvalet and Bengio, 2004). He et al. (2020) extended this approach to sequence generation NLP tasks, highlighting that biased sampling and noise perturbation are key factors for the success of self-training in these tasks. Pace et al. (2024) extended the paradigm of self-training to the alignment problem, improving the robustness of reward models by allowing them to iteratively learn from the highest-scoring and lowest-scoring answers in the candidate pool generated for a query.
@@
To solidify such enhancement, Self-Improve (Huang et al., 2023a) views CoT with Self-Consistency as a policy improvement operator, substantially improving the reasoning capabilities and potential of LLMs through iterative learning of the reasoning paths obtained from Self-Consistency. Zhang and Parkes (2023) demonstrate that LMs could self-improve on large number addition problems by curriculum “distilling” the CoT answers to the direct answers without explicit reasoning. Quiet-STAR (Zelikman et al., 2024) considers the advantage of the influence of rollout rationale on the probabilities of subsequent tokens as a feedback signal, encouraging the model to generate a more helpful implicit thought process using reinforcement learning techniques. Li and Qiu (2023) show that memory mechanisms can also be used for consistency-based self-improvement instead of parameter learning.
@@ 3.1.2 H2: LLMs can judge, critic, refine and more
Judge refers to determining the quality of model responses. The judgment standards are typically incorporated into instructions as principles or guidelines, enabling regulators to oversee LLM behavior in a more scalable manner (Bai et al., 2022b; Yuan et al., 2024c) compared to heavy reliance on human annotators for feedback (Bai et al., 2022a). This approach allows for timely regulation, aiding in the flexible and controlled alignment process of language models, which can help prevent issues like reward hacking during iterative training (Sun et al., 2023c) and facilitate on-policy reinforcement learning training (Guo et al., 2024a).
Self-judging can manifest in two primary forms: 1) Differentiating the relative quality of two responses (AI Feedback (Bai et al., 2022b)), resulting in evaluation outcomes represented as a partial order. For instance, Tan et al. (2023) employ prompts to compare which answer better adheres to the HHH principles. They then distill the chosen option back into the model to further enhance its judging capability. Bai et al. (2022b) prompted the model to select superior responses based on sampled principles and subsequently employed a preference process to achieve a Pareto improvement of the model. 2) Providing an absolute score for a response (LLM-as-a-judge (Zheng et al., 2024b)), with the evaluation results in scalar form. Yao et al. (2023b), Besta et al. (2024) and Xie et al. (2023) introduce real-time evaluation modules for thought states during reasoning process. These modules serves as a prior during the search process, assisting the model in exploring the action space for questions requiring deliberate thinking. Similarly, RAIN (Li et al., 2024h) utilizes a binary scoring prompt for self-evaluating whether the generation is likely to be harmful, thereby enhancing response safety combined with an inference-time tree search. Yuan et al. (2024c) employ a five-point judge prompt to score the model’s instruction-response outputs, then converted the scores into a partial order for iterative training using DPO.
@@
Critique refers to generating modification suggestions. By leveraging LLM itself for critique, the suggestions can address errors and deficiencies, such as mistakes in summarization (Saunders et al., 2022), machine translation (Fernandes et al., 2023), math reasoning (Lin et al., 2024b), decision-making and programming tasks (Saunders et al., 2022; Shinn et al., 2023). The suggestions can also pertain to abstract value criteria, such as the HHH related principles (Chen et al., 2024e; Bai et al., 2022b).
Refine refers to the ability that LLMs can improve the given text. Most of the work on self-refine is based on natural language reasons provided by the critic module to modify the response (Bai et al., 2022b; Tan et al., 2023; Madaan et al., 2023; Shinn et al., 2023). Some research also demonstrates the possibility of making modifications directly based on scalar rewards (Shinn et al., 2023). The less informative critic can be more challenging for LLMs since they must complete more information by themselves through reasoning. Another line of work uses LLMs to refine the prompts themselves (Fernando et al., 2023; Yang et al., 2023a).
@@
    Conducting causal studies on the formation process of models’ judging, critiquing, and refining abilities, and investigating what forms of pre-training and fine-tuning data can influence these abilities. This helps to design specific pipelines to targeted enhance these capabilities (Wang et al., 2024g), and even train specialized models like CriticGPT (McAleese et al., 2024).
@@ 3.1.3 H3: LLMs can effectively learn in-context
However, prepending few-shot exemplars in the context above can make inference inefficient (Gim et al., 2023) and interfered with the unrelated queries (Shi et al., 2023). Therefore, self-generated labels obtained from ICL could be directly used as pseudo labels, and distilled back into the LLMs only paired with the query. This paradigm is known as Context Distillation (Askell et al., 2021; Snell et al., 2022). For instance, in the alignment process of Llama-2 (Touvron et al., 2023), Context Distillation is used to alleviate the problem of long-term dependencies of system prompts. For Llama-3 (Dubey et al., 2024), context distillation is utilized to steer the model towards generating more readable and well-documented code. In Dromedary (Sun et al., 2023d), the base language model is transformed into a safe and helpful aligned model with minimal annotations by directly training on samples acquired from multiple ICL processes. Padmanabhan et al. (2023) demonstrate that Context Distillation can also be used to inject new knowledge to models by learning continuations from entity definitions. Furthermore, Yang et al. (2024b) illustrate the effectiveness of distilling the preference pair generated by contrastive in-context constraints back into the model.
@@ 3.2.2 H5: Self-play
As discussed in H2, the judge and critic model is commonly considered a type of discriminator. For example, Yuan et al. (2024c) utilized rewards from LLM-as-a-Judge to identify high-quality and low-quality responses from the generator, optimizing the generator towards the higher quality ones. However, the adversarial setting between the discriminator and the generator is limited because the only assumption is that the discriminator’s ability can be improved with general helpful training. The discriminator remains almost static (the prompts are unchanged) during training, making it possible for the generator to be over-optimized against the discriminator, leading to reward hacking. Effectively improving the judge and critic module alongside the generator is thus a crucial problem. One reasonable approach is to formulate the training process as an adversarial game (Cheng et al., 2023), where the policy and reward model are updated alternatively via a min-max loss. Another way to introduce a more adversarial setting is to optimize the gaming problem at inference time, as demonstrated in the Consensus Game by Jacob et al. (2024). It employs the piKL no-regret learning algorithm to iteratively update the strategies of both the generator and discriminator, converging to a Nash equilibrium. This equilibrium strategy is then used to rank candidate responses, prioritizing those agreed upon by both players.
@@
The debate paradigm (Irving et al., 2018) is largely inspired by factored cognition and AlphaGo (Silver et al., 2018). In the learning algorithm of AlphaGo, three distinct components are integrated: a player, a counterpart (itself), and a value model that evaluates the win rate associated with each board state. By employing Monte Carlo Tree Search (MCTS), the algorithm conducts rollouts, which are simulated self-play trajectories that extend until a game’s conclusion. These rollouts enhance the accuracy of the value estimates through backward updates based on the outcomes, concurrently refining the policy by capitalizing on strategies that have previously led to victories.
@@
The game of Go shares similarities with tackling the scalable oversight problem using natural language debate. In the beginning or middle of a Go game, even experienced experts may find it difficult to judge which side has a higher probability of winning, just as humans have a small probability of correctly judging problems that exceed human knowledge levels. However, as the game nears its end, the outcome usually becomes clear that even a non-expert judge can confidently evaluate the board generated by the Go masters. For a debate competition, the winner can typically be summarized by the judges. This analogy illustrates a crucial point: through the skillful introduction of adversarial processes, the burden of oversight for complex problems can be significantly reduced.
@@
This provide a possible oversight solution to build trustworthy superhuman AI systems. Irving et al. (2018) show that honesty is better strategy than lie in debate paradigm through proof-of-concept experiments. As an extension of this, Brown-Cohen et al. (2023) propose a new set of debate protocols, wherein the honest strategy can always succeed through a simulation involving only a polynomial number of steps. Khan et al. (2024) conduct a thorough empirical study on the feasibility of implementing the debate paradigm on LLMs: it was found that the debate paradigm can significantly enhance truthfulness, and more persuasive (Anthropic, 2024a) debaters lead to more truthful outcomes. Furthermore, Kirchner et al. (2024) demonstrate that a prover-verifier game employing a powerful honest prover, a potent malicious prover, and a comparatively weaker verifier can generate more readable outputs, thus enabling more effective human oversight of the stronger model.
@@
Apart from the classic natural language debate, a growing body of research has explored the implementation of the debate paradigm across diverse game scenarios to enhance specific model features. A representative arena is the bargaining task (Nash et al., 1950). Fu et al. (2023a) focus on zero-sum variants of bargaining, where the balloon seller aims to sell at a higher price while the buyer seeks a lower price. They observed significant variations in bargaining capabilities among different LLMs and their capacity to learn from play experiences and feedback. Cheng et al. (2024) implement the adversarial language game adversarial taboo (Yao et al., 2021), where an attacker and a defender engage in a conversation centered around a target word visible only to the attacker. The attacker subtly induces the defender to unconsciously utter the target word, while the defender tries to avoid doing so and guess the word from the context. Both players acquire basic gaming skills through imitation learning from a teacher LLM and then refine their strategies through self-play. Interestingly, less capable player LLMs not only improve their win rates in this specific game but also enhance their general reasoning abilities. Ma et al. (2023a) introduce the red-teaming game, a more intricate adversarial team game where LLMs are initialized as a joint set of red-teaming policies to prompt the target LLM to produce harmful content. They propose a solver to ensure the final meta-strategy approximates a Nash equilibrium within a certain $\epsilon$ margin. Zheng et al. (2024c) suggests addressing the alignment problem by allowing an attacker to prompt a defender LLM to generate answers that might result in low rewards, while the defender tries to maximize the rewards of these prompts. The solution of this game is considered an iterative min-max optimization process with constraints.